# KG-Infused RAG: Augmenting Corpus-Based RAG with External Knowledge Graphs

### v1

@@ 3.5 KG-Based Query Expansion
The goal of this stage is to generate an expanded query $q^{\prime}$ that complements the original query $q$ by incorporating structured knowledge retrieved from the KG. The expansion broadens retrieval coverage and enhances the relevance of corpus evidence.
We prompt the LLM with both the original query $q$ and the KG subgraph summary $\mathcal{S}_{\mathcal{G},q}^{act}$ constructed via spreading activation. Conditioned on the prompt, the model performs associative reasoning to transform the query—either simplifying it based on retrieved facts (e.g., replacing intermediate entities), or extending it with new, KG-grounded content inferred through LLM knowledge. Formally,
$q^{\prime}=\text{LLM}_{\text{prompt}}\left(q,\mathcal{S}_{\mathcal{G},{q}}^{act}\right)$ (7)
We then perform dual-query retrieval over the corpus using both $q$ and $q^{\prime}$, and merge the results:
@@ 3.6 KG-Augmented Answer Generation
This stage integrates corpus and KG evidence to generate the final answer. We enhance retrieved passages with structured facts to support more informed and accurate reasoning.
@@
We first prompt the LLM to summarize the retrieved passages $\mathcal{D}_{c,(q,q^{\prime})}$, producing a query-focused passage note $\mathcal{S}_{\mathcal{D},q}$ that distills key information relevant to answering $q$:
$\mathcal{S}_{\mathcal{D},q}=\text{LLM}_{\text{prompt}}\left(q,\mathcal{D}_{c,(q,q^{\prime})}\right)$ (9)
@@
$\mathcal{S}_{q}^{\text{final}}=\text{LLM}_{\text{prompt}}\left(\mathcal{S}_{\mathcal{D},q},\mathcal{S}_{\mathcal{G},q}^{act}\right)$ (10)
@@
Finally, the LLM generates the answer conditioned on the original query $q$ and the enriched note $\mathcal{S}_{q}^{\text{final}}$:
$\text{answer}=\text{LLM}_{\text{prompt}}\left(q,\mathcal{S}_{q}^{\text{final}}\right)$ (11)
This design enables the model to generate answers grounded not only in textual evidence but also in structured KG facts, facilitating more comprehensive, accurate, and interpretable reasoning.
3.7 Training: Data Construction and DPO Optimization
To improve the instruction-following ability at key stages of KG-Infused RAG’s pipeline, we construct a simple preference-based training dataset by sampling a single key stage from the pipeline and train the generator used in multiple stages using Direct Preference Optimization (DPO) [23].
DPO Data Construction.
We use an advanced LLM, GPT-4o-mini, to construct a training dataset for DPO based solely on outputs from a single stage—KG-guided knowledge augmentation. Specifically, we sample examples from the 2WikiMultiHopQA [12] training set and collect intermediate outputs from this stage to form the dataset $D_{\text{KA}}$. For each input $x$, we generate multiple candidate outputs using diverse decoding strategies, and prompt GPT-4o-mini to identify the best and worst outputs. We construct preference-labeled triples $(x,y^{-},y^{+})$. The resulting dataset is denoted as:
DPO Training.
To train the model to prefer higher-quality knowledge-augmented outputs, we apply the DPO objective. Given a preference-labeled triple $(x,y^{-},y^{+})$ from $D_{\text{DPO}}$, the DPO loss encourages the policy model $\pi_{\theta}$ to assign higher likelihood to the preferred output $y^{+}$, while maintaining consistency with a fixed reference model $\pi_{\text{ref}}$. The training objective is defined as:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\text{ref}}(y^{+}|x)}-\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{\text{ref}}(y^{-}|x)}\right)\right)\right]$ (13)
where $\pi_{\theta}$ denotes the trainable model, and $\pi_{\text{ref}}$ is a fixed reference model. Additional training details are provided in Appendix E.
@@ 4.1 Datasets & Evaluation
To evaluate our method in complex QA scenarios, we conduct experiments on four challenging multi-hop QA datasets: HotpotQA [31], 2WikiMultihopQA (2WikiMQA) [12], MuSiQue [24], and Bamboogle [22]. For HotpotQA, 2WikiMQA, and MuSiQue, we use the subsets provided by Trivedi et al. [25]. These datasets require models to perform complex multi-hop reasoning across multiple pieces of evidence, making them ideal for evaluating our method’s ability to integrate and reason over knowledge. We report F1 Score (F1), Exact Match (EM), Accuracy (Acc), and their average (Avg) as evaluation metrics. The summary of datasets and evaluation is in Appendix C.1.
@@
To assess the generalizability of our method to simpler QA settings, we evaluate it on the StrategyQA dataset [9], a commonsense QA task requiring binary Yes/No answers. Unlike multi-hop QA tasks that demand explicit retrieval and integration of multiple evidence pieces, StrategyQA emphasizes implicit reasoning based on implicit commonsense knowledge. From the RAG perspective, it poses a lower retrieval challenge. We randomly sample 500 examples from the test set for evaluation, with accuracy as the evaluation metric.
@@ 4.2 Baselines
We compare our method against three standard baselines: (1) No-Retrieval (NoR) directly feeds the input query into an LLM without retrieving any external knowledge. (2) Vanilla RAG retrieves the top-$k_{p}$ passages from the corpus based on the raw query $q$. (3) Vanilla Query Expansion (Vanilla QE) prompts an LLM to generate an expanded query $q^{\prime}$ without incorporating KG facts. Both the original and expanded queries are used to retrieve top-$k_{p}/2$ passages each. For a fair comparison, both Vanilla RAG and Vanilla QE adopt our passage note construction step to form passage notes $\mathcal{S}_{\mathcal{D},q}$, which are then used by the LLM to generate answers.
@@
To evaluate the generalizability of our method as a plug-and-play module, we integrate it into two strong adaptive RAG methods: Self-RAG [2] and Self-Ask [22].
For the Self-RAG baseline, since all evaluation datasets are short-form generation tasks, we follow the official setting where multiple passages are retrieved only once for generation. To apply KG-Infused RAG as a plug-in to Self-RAG, we first perform KG-Based QE after KG-Guided Spreading Activation to retrieve new passages, replacing the originals. We then apply KG-Aug Gen to each retrieved passage individually. The subsequent generation process remains unchanged.
The details of the Self-Ask baseline and its combination with KG-Infused RAG are in Appendix F.1.
@@ 4.3 Resources and Model Components
To ensure consistency, we use the same KG and text corpus across all datasets and baselines. For the KG $\mathcal{G}$, we utilize Wikidata5M-KG, an open-domain KG derived from Wikidata entities aligned with Wikipedia. Preprocessing details are provided in Section 3.3. For the corpus $\mathcal{D}_{c}$, we adopt Wikipedia-2018, a large-scale open-domain corpus containing 21,015,324 passages [^1] https://github.com/facebookresearch/DPR.
@@
We adopt Contriever-MS MARCO [13] as the retriever for both entity and passage retrieval across all experiments. For generation, Qwen2.5-7B [30] and LLaMA3.1-8B [10] serve as base generators in our method. For Self-RAG, we employ selfrag-llama2-7b [^2] https://huggingface.co/selfrag/selfrag_llama2_7b as the generator, which is additionally fine-tuned following the original Self-RAG paper. For the Self-Ask baseline, Qwen2.5-7B is selected for its superior instruction-following capability in this setting. The DPO training dataset is constructed using gpt-4o-mini-0718 from the GPT-4o family [1].
4.3.1 Method Configuration
Passage Retrieval.
For all baselines and our method, the default number of retrieved passages $k_{p}$ is set to 6. In both Vanilla-QE and KG-Infused RAG, the original query $q$ and the expanded query $q^{\prime}$ each retrieve $k_{p}/2=3$ passages from the corpus. Further details on passage retrieval configuration can be found in Appendix C.2.
Initial Entities Retrieval.
For KG-Infused RAG, we begin by retrieving the top $k_{e}$ entities from KG, where $k_{e}$ is set to 3 by default.
Other Implementation Details.
We set the maximum number of activation rounds to 6. Because each round may extract numerous triples, potentially introducing noise that degrades model performance, we introduce two hyperparameters to control the expansion scope:
  • MAX_ENTITIES_PER_ROUND: Limits the number of new entities $\{e^{\prime}\}$ introduced in each activation round, preventing excessive expansion and potential noise.
  • MAX_TRIPLES_PER_ENTITY: Restricts the number of triples retrieved per entity, reducing noise from entities with an excessive number of triples (e.g., countries or regions).
Table 1: Overall results (%) on multiple datasets. Bold numbers denote the best-performed results. “w/ $n$-Round” denotes KG-Infused RAG with $n$ activation rounds.
Method Multi-hop Short-form
@@ 4.3.1 Method Configuration
Qwen2.5-7B
NoR 19.8 25.8 18.0 21.2 23.4 26.8 22.0 24.1 4.0 11.6 3.4 6.3 10.4 16.6 8.0 11.7 64.4
Vanilla RAG 35.0 41.9 31.0 36.0 30.6 33.5 27.2 30.4 5.2 11.5 3.6 6.8 22.4 30.1 21.6 24.7 68.6
Vanilla QE 35.0 41.8 30.8 35.9 28.6 31.5 25.4 28.5 5.4 11.1 2.2 6.2 20.8 28.6 18.4 22.6 72.2
KG-Infused RAG (w/ 2-Round) 39.0 44.7 34.4 39.4 44.8 44.5 36.0 41.8 10.8 16.7 7.6 11.7 26.4 34.1 24.0 28.2 68.2
                +DPO (w/ 2-Round) 37.2 43.9 31.0 37.4 45.6 45.8 35.8 42.4 12.8 19.2 8.6 13.5 36.8 41.8 31.2 36.6 72.8
$\Delta_{\text{KG-Infused RAG$\rightarrow$Vanilla RAG}}$ 4.0$\uparrow$ 2.8$\uparrow$ 3.4$\uparrow$ 3.4$\uparrow$ 14.2$\uparrow$ 11.0$\uparrow$ 8.8$\uparrow$ 11.4$\uparrow$ 5.6$\uparrow$ 5.2$\uparrow$ 4.0$\uparrow$ 4.9$\uparrow$ 4.0$\uparrow$ 4.0$\uparrow$ 2.4$\uparrow$ 3.5$\uparrow$ -0.4$\downarrow$
$\Delta_{\text{KG-Infused RAG+DPO$\rightarrow$Vanilla RAG}}$ 2.2$\uparrow$ 2.0$\uparrow$ 0.0$\uparrow$ 1.4$\uparrow$ 15.0$\uparrow$ 12.3$\uparrow$ 8.6$\uparrow$ 12.0$\uparrow$ 7.6$\uparrow$ 7.7$\uparrow$ 5.0$\uparrow$ 6.7$\uparrow$ 14.4$\uparrow$ 11.7$\uparrow$ 9.6$\uparrow$ 11.9$\uparrow$ 4.2$\uparrow$
LLaMA3.1-8B
NoR 22.2 27.0 20.6 23.3 28.0 29.3 24.4 27.2 3.6 8.8 3.0 5.1 11.2 16.4 8.0 11.9 68.4
Vanilla RAG 32.2 39.0 28.8 33.4 30.6 31.6 25.0 29.1 3.6 9.6 2.0 5.1 20.0 24.8 18.4 21.1 66.2
Vanilla QE 32.6 38.8 28.8 33.4 26.6 28.6 21.6 25.6 5.2 11.0 3.4 6.5 18.4 23.3 15.2 19.0 68.0
KG-Infused RAG (w/ 2-Round) 34.4 39.8 30.2 34.8 37.0 36.0 27.4 33.5 12.0 16.9 7.4 12.1 24.0 28.5 22.4 25.0 67.0
                +DPO (w/ 2-Round) 36.6 43.8 31.0 37.1 45.0 46.8 37.0 42.9 12.2 18.5 8.8 13.2 31.2 33.6 24.8 29.9 69.4
$\Delta_{\text{KG-Infused RAG$\rightarrow$Vanilla RAG}}$ 2.2$\uparrow$ 0.8$\uparrow$ 1.4$\uparrow$ 1.4$\uparrow$ 6.4$\uparrow$ 4.4$\uparrow$ 2.4$\uparrow$ 4.4$\uparrow$ 8.4$\uparrow$ 7.3$\uparrow$ 5.4$\uparrow$ 7.0$\uparrow$ 4.0$\uparrow$ 3.7$\uparrow$ 4.0$\uparrow$ 3.9$\uparrow$ 0.8$\uparrow$
$\Delta_{\text{KG-Infused RAG+DPO$\rightarrow$Vanilla RAG}}$ 4.4$\uparrow$ 4.8$\uparrow$ 2.2$\uparrow$ 3.7$\uparrow$ 14.4$\uparrow$ 15.2$\uparrow$ 12.0$\uparrow$ 13.8$\uparrow$ 8.6$\uparrow$ 8.9$\uparrow$ 6.8$\uparrow$ 8.1$\uparrow$ 11.2$\uparrow$ 8.8$\uparrow$ 6.4$\uparrow$ 8.8$\uparrow$ 3.2$\uparrow$

### v2

@@ 3.4 KG-Based Query Expansion
The goal of this stage is to generate an expanded query $q^{\prime}$ which complements the original query $q$ by incorporating structured knowledge retrieved from the KG. The expansion broadens the retrieval coverage and enhances the relevance of corpus evidence.
We prompt the LLM with both the original query $q$ and the KG subgraph summary $\mathcal{S}_{\mathcal{G},q}^{act}$ constructed via spreading activation. Conditioned on the prompt, the model performs associative reasoning to transform the query—either simplifying it based on retrieved facts (e.g., replacing intermediate entities), or extending it with new, KG-grounded content inferred from LLM knowledge. Formally,
$q^{\prime}\sim\text{LLM}\left(\text{Instruct}_{\text{QE}},q\parallel\mathcal{S}_{\mathcal{G},{q}}^{act}\right)$ (6)
Where $\text{Instruct}_{\text{QE}}$ is the prompt template used for query expansion. We then perform dual-query retrieval over the corpus using both $q$ and $q^{\prime}$, and merge the results:
@@ 3.5 KG-Augmented Answer Generation
This stage integrates both corpus and KG evidence to generate the final answer. By augmenting the retrieved passages with structured facts, we facilitate more informed and accurate reasoning.
@@
We first prompt the LLM to summarize the retrieved passages $\mathcal{D}_{c,(q,q^{\prime})}$, generating a query-focused passage note $\mathcal{S}_{\mathcal{D},q}$ that distills key information relevant to answering $q$:
$\mathcal{S}_{\mathcal{D},q}\sim\text{LLM}\left(\text{Instruct}_{\text{Note}},q\parallel\mathcal{D}_{c,(q,q^{\prime})}\right)$ (8)
where $\text{Instruct}_{\text{Note}}$ is the prompt template used for passage note construction.
@@
$\mathcal{S}_{q}^{\text{final}}\sim\text{LLM}\left(\text{Instruct}_{\text{KA}},\mathcal{S}_{\mathcal{D},q}\parallel\mathcal{S}_{\mathcal{G},q}^{act}\right)$ (9)
where $\text{Instruct}_{\text{KA}}$ is the prompt template used for knowledge augmentation.
@@
Finally, the LLM generates the answer conditioned on the original query $q$ and the enriched note $\mathcal{S}_{q}^{\text{final}}$, as follows:
$\text{answer}\sim\text{LLM}\left(\text{Instruct}_{\text{Ans}},q\parallel\mathcal{S}_{q}^{\text{final}}\right)$ (10)
where $\text{Instruct}_{\text{Ans}}$ is the prompt template used for answer generation. This design enables the model to generate answers grounded not only in textual evidence but also in structured KG facts, facilitating more comprehensive, accurate, and interpretable reasoning.
3.6 Direct Preference Optimization for Knowledge Augmentation
Inspired by previous works (Li et al., 2025; Wang et al., 2024), we enhance the instruction-following capability of the model at the KG-Guided KA stage of KG-Infused RAG’s pipeline by using Direct Preference Optimization (DPO) (Rafailov et al., 2023), which aligns with better response preferences, thereby enabling better integration of knowledge retrieved from both the KG and the corpus.
We construct a simple, preference-based training dataset $D_{\text{KA}}$ by sampling the response from the KG-Guided KA stage. The training objective is defined as:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\text{KI}}^{\theta}(y^{+}|x)}{\pi_{\text{KI}}^{\text{ref}}(y^{+}|x)}-\log\frac{\pi_{\text{KI}}^{\theta}(y^{-}|x)}{\pi_{\text{KI}}^{\text{ref}}(y^{-}|x)}\right)\right)\right]$ (11)
where $\pi_{\text{KI}}^{\theta}$ denotes the trainable model, and $\pi_{\text{KI}}^{\text{ref}}$ is a fixed reference model during training. Training data construction and additional training details are provided in Appendix D.
@@ 4.1 Datasets & Evaluation
To evaluate our method in complex QA scenarios, we conduct experiments on four challenging multi-hop QA datasets: HotpotQA (Yang et al., 2018), 2WikiMultihopQA (2WikiMQA) (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and Bamboogle (Press et al., 2023). For HotpotQA, 2WikiMQA, and MuSiQue, we use the subsets provided by Trivedi et al. (2023). These datasets require models to perform complex multi-hop reasoning across multiple pieces of evidence, making them ideal for evaluating our method’s ability to integrate and reason over knowledge. We report F1 Score (F1), Exact Match (EM), Accuracy (Acc), and their average (Avg) as evaluation metrics. The summary of datasets and evaluation is in Appendix B.1.
@@
To assess the generalizability of our method to simpler QA scenarios, we evaluate it on the StrategyQA dataset (Geva et al., 2021), a commonsense QA task requiring binary Yes/No answers. Unlike multi-hop QA tasks that demand explicit retrieval and integration of multiple evidence pieces, StrategyQA emphasizes implicit reasoning based on implicit commonsense knowledge. From the RAG perspective, it poses a lower retrieval challenge. We randomly sample 500 examples from the test set for evaluation, with accuracy as the evaluation metric.
@@ 4.2 Baselines
We compare our method against four types of baselines: (1) No-Retrieval (NoR), which directly feeds the query into an LLM to generate the answer. (2) Vanilla RAG, which retrieves the top-$k_{p}$ passages from the corpus using the raw query $q$, and generate the answer. (3) Vanilla Query Expansion (Vanilla QE), which first prompts an LLM to generate an expanded query $q^{\prime}$ without leveraging KG facts, then both the raw and the expanded queries are used to retrieve top-$k_{p}/2$ passages each, and generate the answer. (4) KG-Based methods, which first use LLMs to construct an ad-hoc KG, then perform retrieval to generate the answer. For Vanilla RAG and Vanilla QE, we adopt our passage note construction step to ensure a fair comparison with KG-Infused RAG. For KG-Based methods, we select three representative baselines, including GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2024), and KG2RAG (Zhu et al., 2025).
@@
To evaluate the generalizability of our method as a plug-and-play module, we integrate it into three representative adaptive RAG methods: Self-RAG (Asai et al., 2024), Self-Ask (Press et al., 2023), and DeepNote (Wang et al., 2024). Specifically, we incorporate the KG-Guided Spreading Activation module into these baselines, enabling them to retrieve semantically related concepts from the KG, and further employ KG-Based QE and KG-Aug Gen to enhance their performance. Detailed integration procedures for each baseline are provided in Appendix E.
@@ 4.3 Resources and Model Components
For the KG $\mathcal{G}$ used in KG-Infused RAG, we adopt Wikidata5M-KG, an open-domain KG derived from Wikidata and Wikipedia. To ensure fair comparisons, we use the same corpus $\mathcal{D}_{c}$ within each group of baselines: (1) for non-KG-based baselines, we use the large-scale Wikipedia-2018 corpus containing 21,015,324 passages [^1] https://github.com/facebookresearch/DPR; (2) for KG-based baselines, given the high cost of KG construction, we employ a smaller corpus by concatenating the HotpotQA, 2WikiMQA, and MuSiQue corpora provided in HippoRAG2 (Gutiérrez et al., 2024), resulting in 27,572 passages.
@@
We adopt Contriever-MS MARCO (Izacard et al., 2022) as the retriever for both entity and passage retrieval in all experiments, except that for certain KG-based baselines, we follow their official implementations. For generation, we employ a diverse set of language models spanning parameter scales, covering both open-source and proprietary settings. Specifically, we evaluate with the Qwen2.5 series (Yang et al., 2024) (3B, 7B, 14B), the LLaMA3.1 series (Grattafiori et al., 2024) (8B, 70B), and the GPT-4o-mini (Hurst et al., 2024).
4.4 Implementation Details.
Knowledge Retrieval.
For all non-KG-based baselines and our method, the default number of retrieved passages $k_{p}$ is set to 6. In both Vanilla-QE and KG-Infused RAG, the original query $q$ and the expanded query $q^{\prime}$ each retrieve $k_{p}/2=3$ passages from the corpus. For GraphRAG and LightRAG, we retain their default top-$k$ retrieval configurations specified in their original implementations. A comprehensive summary of the retrieval configurations for all methods is provided in Appendix B.2.
KG-Infused RAG Implementation Details.
Table 1: Overall results (%) on five datasets. Bold numbers denote the best-performed results. We report KG-Infused RAG with two activation rounds (“w/ 2-Round”), where “w/ n-Round” denotes KG-Infused RAG equipped with $n$ rounds of activation.
Method & LLMs Multi-hop Commonsense
@@
No-Retrieval (NoR)
Qwen2.5-3B-Instruct 15.4 21.1 14.8 17.1 25.0 28.3 24.4 25.9 1.2 6.8 0.6 2.9 6.4 9.0 3.2 6.2 61.2
Qwen2.5-7B-Instruct 19.8 25.8 18.0 21.2 23.4 26.8 22.0 24.1 4.0 11.6 3.4 6.3 10.4 16.6 8.0 11.7 64.4
LLaMA3.1-8B-Instruct 22.2 27.0 20.6 23.3 28.0 29.3 24.4 27.2 3.6 8.8 3.0 5.1 11.2 16.4 8.0 11.9 68.4
Qwen2.5-14B-Instruct 22.2 28.8 21.8 24.3 28.6 31.2 27.6 29.1 5.6 12.0 4.4 7.3 13.6 20.8 12.0 15.5 64.4
GPT4o-mini 31.0 38.9 29.4 33.1 29.4 32.6 25.6 29.2 7.4 15.1 5.4 9.3 19.2 28.1 18.4 21.9 74.6
LLaMA3.1-70B-Instruct 31.8 40.5 30.4 34.2 32.6 36.5 30.4 33.2 8.0 13.8 6.4 9.4 32.0 41.5 28.8 34.1 72.0
Vanilla RAG
Qwen2.5-3B-Instruct 28.2 34.8 26.0 29.7 29.0 32.3 27.2 29.5 3.8 9.6 3.2 5.5 12.0 19.6 9.6 13.7 62.6
Qwen2.5-7B-Instruct 35.0 41.9 31.0 36.0 30.6 33.5 27.2 30.4 5.2 11.5 3.6 6.8 22.4 30.1 21.6 24.7 68.6
LLaMA3.1-8B-Instruct 32.2 39.0 28.8 33.4 30.6 31.6 25.0 29.1 3.6 9.6 2.0 5.1 20.0 24.8 18.4 21.1 66.2
Qwen2.5-14B-Instruct 35.2 42.6 32.4 36.7 32.6 30.1 23.6 28.8 6.8 11.8 4.4 7.7 26.4 34.8 25.6 28.9 70.6
GPT4o-mini 43.0 52.4 39.6 45.0 33.6 36.0 28.4 32.7 9.8 18.1 8.0 12.0 32.0 41.1 30.4 34.5 74.4
LLaMA3.1-70B-Instruct 37.4 46.0 34.8 39.4 28.6 29.6 24.6 27.6 8.2 13.0 7.0 9.4 29.6 35.8 29.6 31.7 71.0
Vanilla Query Expansion (Vanilla QE)
Qwen2.5-3B-Instruct 27.6 34.6 25.4 29.2 28.6 32.7 27.2 29.5 4.2 9.4 3.6 5.7 15.2 23.8 13.6 17.5 61.4
Qwen2.5-7B-Instruct 35.0 41.8 30.8 35.9 28.6 31.5 25.4 28.5 5.4 11.1 2.2 6.2 20.8 28.6 18.4 22.6 72.2
LLaMA3.1-8B-Instruct 32.6 38.8 28.8 33.4 26.6 28.6 21.6 25.6 5.2 11.0 3.4 6.5 18.4 23.3 15.2 19.0 68.0
Qwen2.5-14B-Instruct 35.0 42.4 32.0 36.5 32.6 31.1 25.2 29.6 7.6 13.3 5.0 8.6 29.6 35.4 29.6 31.5 72.4
GPT4o-mini 43.0 52.2 38.8 44.7 34.4 37.8 30.2 34.1 11.0 18.6 9.2 12.9 32.8 40.9 31.2 35.0 74.6
LLaMA3.1-70B-Instruct 38.4 45.9 36.0 40.1 26.8 28.0 23.4 26.1 11.2 15.2 8.6 11.7 33.6 39.4 31.2 34.7 72.6
KG-Infused RAG (ours)
KG-Infused RAG (Qwen2.5-3B-Instruct) 28.0 34.6 25.4 29.3 39.0 41.8 34.0 38.3 8.6 14.0 7.0 9.9 16.0 23.9 14.4 18.1 61.4
KG-Infused RAG (Qwen2.5-7B-Instruct) 39.0 44.7 34.4 39.4 44.8 44.5 36.0 41.8 10.8 16.7 7.6 11.7 26.4 34.1 24.0 28.2 68.2
       +DPO 37.2 43.9 31.0 37.4 45.6 45.8 35.8 42.4 12.8 19.2 8.6 13.5 36.8 41.8 31.2 36.6 72.8
KG-Infused RAG (LLaMA3.1-8B-Instruct) 34.4 39.8 30.2 34.8 37.0 36.0 27.4 33.5 12.0 16.9 7.4 12.1 24.0 28.5 22.4 25.0 67.0
       +DPO 36.6 43.8 31.0 37.1 45.0 46.8 37.0 42.9 12.2 18.5 8.8 13.2 31.2 33.6 24.8 29.9 69.4
[…]
KG-Infused RAG (Qwen2.5-14B-Instruct) 38.0 45.6 35.4 39.7 47.4 44.9 35.6 42.6 14.0 19.0 9.4 14.1 31.2 42.6 29.6 34.5 75.2
KG-Infused RAG (GPT4o-mini) 44.0 53.5 39.8 45.8 47.6 47.5 38.0 44.4 15.0 21.9 10.8 15.9 38.4 43.3 36.0 39.2 75.4
KG-Infused RAG (LLaMA3.1-70B-Instruct) 41.6 50.1 38.2 43.3 49.0 49.2 40.8 46.3 17.0 22.7 13.0 17.6 48.8 54.2 45.6 49.5 75.0
[…]
Table 2: Results (%) of KG-based methods. “local” and “global” indicate different search modes used in these methods.
Method HotpotQA 2WikiMQA MuSiQue Bamboogle
Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg
GraphRAG (local) (GPT-4o-mini) 15.0 15.3 10.4 13.6 13.8 8.7 5.8 9.4 5.6 7.0 3.0 5.2 12.8 12.2 9.6 11.5
GraphRAG (global) (GPT-4o-mini) 32.4 28.9 20.2 27.2 31.2 19.2 13.2 21.2 10.8 9.4 3.0 7.7 44.0 34.0 28.8 35.6
LightRAG (local) (GPT-4o-mini) 28.0 30.0 20.6 26.2 34.2 30.8 24.0 29.7 18.6 19.5 9.4 15.8 32.0 27.6 16.0 25.2
LightRAG (global) (GPT-4o-mini) 27.6 27.3 18.2 24.4 36.0 24.7 17.6 26.1 18.4 16.2 6.4 13.7 30.4 23.1 13.6 22.4
KG2RAG (LLaMA3.1-8B-Instruct) 21.4 25.4 18.0 21.6 29.4 19.3 11.4 20.0 13.8 18.7 10.2 14.2 9.6 13.3 8.0 10.3
KG-Infused RAG (GPT-4o-mini) 35.2 43.8 32.8 37.3 40.8 40.5 31.4 37.6 24.4 30.8 18.8 24.7 40.8 44.7 37.6 41.0
Retrieval is initialized by selecting the top $k_{e}=3$ entities from the KG. The maximum number of activation rounds is set to six. To control expansion and mitigate noise introduced by overly large subgraphs, we apply two constraints: (1) MAX_ENTITIES_PER_ROUND: Limits the number of new entities $\{e^{\prime}\}$ introduced in each activation round, preventing excessive expansion and potential noise. (2) MAX_TRIPLES_PER_ENTITY: Restricts the number of triples retrieved per entity, reducing noise from entities with disproportionately large numbers of triples (e.g., countries or regions).